# Reason-RFT: Reinforcement Fine-Tuning for Visual Reasoning

### v1

@@ 4.1 Experimental Details
Table 1: Results on three visual reasoning tasks. The best results among different training paradigms are highlighted in bold, while the second-best results are underlined. “ID” denotes in-domain test data, and “OOD” denotes out-of-domain test data.
@@
Clevr-Math Super-Clevr AVG GeoMath Geometry3k AVG TRANCE TRANCE-L TRANCE-R AVG
ID OOD ID OOD ID OOD
@@
Qwen2.5-VL-3B-Instruct [5] 75.90 39.30 57.60 36.75 37.44 37.09 8.57 8.26 8.31 8.42
Phi-3.5-Vision-4B-Instruct [1] 21.40 15.20 18.30 36.83 50.25 43.54 7.42 2.45 4.02 5.33
Llava-OneVision-7B [35] 69.70 29.10 49.40 77.63 43.66 60.64 10.00 8.33 8.74 9.27
Qwen2.5-VL-7B-Instruct [5] 74.60 35.20 54.90 44.00 45.61 44.80 19.63 13.12 13.42 16.45
InternVL-2.5-8B [6] 93.50 35.30 64.40 63.00 47.32 51.60 7.19 6.62 6.63 6.91
Llama-3.2-11B-Vision [44] 10.30 9.50 9.90 13.75 20.85 17.30 8.22 8.40 9.03 8.47
Pixtral-12B [3] 42.60 22.90 32.75 30.38 36.09 33.23 7.35 5.03 5.22 6.42
@@
Zero-Shot 82.40 32.00 57.20 25.86 20.63 23.25 3.78 4.60 4.67 4.35
$+$ ANS-SFT 96.20 39.20 67.70 51.34 22.50 36.92 77.39 49.24 50.33 58.99
$+$ CoT-SFT 85.50 46.50 66.00 43.05 25.25 34.15 64.37 43.19 42.86 50.14
$+$ Reason-RFT-Zero 98.40 44.80 71.60 47.68 32.50 40.09 42.13 34.07 33.41 33.74
$+$ Reason-RFT 96.80 51.20 74.00 49.03 33.13 41.08 74.61 64.05 64.08 67.58
@@
Zero-Shot 98.60 42.10 70.35 43.30 43.88 43.59 13.53 12.72 12.78 13.01
$+$ ANS-SFT 95.00 33.90 64.45 51.34 25.38 38.36 82.19 54.29 54.83 63.77
$+$ CoT-SFT 87.30 42.40 64.85 50.49 33.00 41.75 81.31 47.90 47.80 59.00
$+$ Reason-RFT-Zero 99.40 53.00 76.20 55.00 54.75 54.88 67.67 57.20 56.15 56.68
$+$ Reason-RFT 95.60 51.00 73.30 59.27 49.25 54.26 79.97 59.36 58.61 65.98
Datasets In this paper, we comprehensively evaluate the visual reasoning capabilities of our method by leveraging six existing datasets, enhanced through subtask categorization, error-prone data filtering, and dataset restructuring. Detailed protocols for data filtering and restructuring are provided in the supplementary materials. Specifically, we define three task categories as follows.
  • Visual Counting is a multimodal reasoning task evaluating the integration of linguistic, visual, and mathematical skills by solving arithmetic problems in 3D block-based scenes, which includes four subtask types, with datasets from CLEVR-Math [38] and Super-CLEVR [37]. Specifically, CLEVR-Math provides 35k training and 1k test samples, while Super-CLEVR serves as an out-of-domain test set with 1k samples to assess generalization.
  • Structure Perception is a visual reasoning task requiring models to perceive and interpret structural information in mathematical geometries, medical imaging, chart layouts, and architectural designs. It includes both choice and non-choice formats, with datasets sourced from Geo170K [12], Math360K [51] (collectively GeoMath), and Geometry3K [40]. GeoMath provides 4.5k training and 820 test samples, while Geometry3K serves as an out-of-domain test set with 800 samples to evaluate cross-domain adaptability.
  • Spatial Transformation is a spatial-visual reasoning task requiring models to infer single-step or multi-step transformation actions by analyzing initial and final visual states for 3D scenes from multiple perspectives (e.g., center, left, right). The dataset TRANCE [23] features four distinct difficulty levels, with 60k training and 6k test samples. Out-of-domain test samples are generated by rendering identical data from left and right viewpoints to rigorously assess robustness to perspective variations.
Evaluation Metrics We use accuracy-rate (Acc) as the primary metric [67]. For numerical answers, correctness is verified by mathematical equivalence to the ground truth. For multiple-choice questions, we perform a string match. For function-type sequences, we use stepwise multi-level evaluation to assess alignment with the correct solution.
Table 2: More detailed results on Super-Clevr. “Encountered” refers to the types of questions the model has previously seen during Clevr-Math training, while “UnEncountered” denotes the complicated types that the model has not encountered (i.e. questions with mixture of addition and subtraction).
Methods Encountered UnEncountered Total
Qwen2VL-2B-Instruct
Zero-Shot 42.67 0.00 32.00
$+$ ANS-SFT 51.07 5.20 39.20
$+$ CoT-SFT 49.73 36.80 46.50
$+$ Reason-RFT-Zero 58.00 5.20 44.80
$+$ Reason-RFT 60.00 28.40 51.20
Qwen2VL-7B-Instruct
Zero-Shot 54.53 4.80 42.10
$+$ ANS-SFT 42.53 8.00 33.90
$+$ CoT-SFT 45.33 33.60 42.40
$+$ Reason-RFT-Zero 63.60 21.20 53.00
$+$ Reason-RFT 56.13 35.60 51.00
Implementation Details We utilize Qwen2-VL-2B and Qwen2-VL-7B [59] as the backbone models for our experiments. Our implementation is built on the open-source frameworks Open-R1 [25] and vLLM [33], ensuring reproducibility and scalability. All experiments were conducted on a cluster of servers, each equipped with 8$\times$×A800 GPUs. For further details, see the supplementary materials.
Training Paradigms and Baselines To assess the performance and generalization of different training strategies, we compare: (1) SFT-based methods—ANS-SFT, which fine-tunes on answer generation, and CoT-SFT, which uses supervised learning with chain-of-thought (CoT) reasoning; and (2) RL-based methods—Reason-RFT-Zero, which applies reinforcement learning without reasoning activation phase, and Reason-RFT, which uses partial CoT data for reasoning activation before RL training. For comprehensive experiments, we use Qwen2-VL-Instruct [59] as the base model, evaluating both 2B and 7B variants to explore the impact of model scale. In addition, we also select the most advanced open-source models  [5, 1, 35, 6, 44, 3] and the proprietary models  [26, 55] available as baselines to evaluate the performance of different training paradigms.
4.2 Results on In-Domain Tasks
To evaluate the In-Domain (ID) performance of Reason-RFT against different training methods and baseline models across visual reasoning tasks, we conducted extensive training and validation on 2B/7B models for three tasks. The results in Tab. 1 show: (1) Visual Counting RL-based methods outperform all open-source and proprietary baseline models as well as SFT-based methods in both 2B and 7B models, with Reason-RFT-Zero in 7B nodel achieving the best performance; (2) Structure Perception RL-based methods surpass SFT-based approaches in 7B model, while ANS-SFT performs best in 2B model. CoT-SFT shows limited improvement, as enforced reasoning supervision inhibits cognitive enhancement. Additionally, Reason-RFT in 7B model outperforms all proprietary and most open-source models except InternVL-2.5-8B [6] and Llava-OneVision-7B [35]; (3) Spatial Transformation SFT-based methods achieve the highest performance, with Reason-RFT achieving comparable performance while totally surpassing all baseline models. However, Reason-RFT-Zero performs poorly, likely due to its difficulty adapting to function-based tasks, whereas Reason-RFT benefits from reasoning activation, reaching a higher performance ceiling. In summary, Reason-RFT achieves performance comparable to SFT-based methods in ID tasks while outperforming all baselines in Visual Counting and Spatial Transformation, and most of baselines in Structure Perception.
4.3 Results on Out-of-Domain Generalization
To validate the Out-of-Domain (OOD) performance of Reason-RFT across different training methods and baseline models in visual reasoning tasks, we conducted comprehensive experiments on 2B/7B models for three tasks. Results in Tab. 1 and Tab. 2 show that: (1) Visual Counting RL-based methods demonstrate superior generalization over SFT-based methods in both 2B and 7B models, with Reason-RFT surpassing ANS-SFT by 12% (2B) and 17% (7B) while outperforming all open-source and proprietary baselines. Notably, compared with purely RL-based method (i.e., Reason-RFT-Zero), CoT-SFT and Reason-RFT can activate model capabilities on unseen complicated questions, achieving strong results despite its poor zero-shot performance before in-domain training, as shown in Tab. 2. This also highlights the limitations of Reason-RFT-Zero due to cognitive constraints, especially in smaller models (i.e. 2B models); (2) Structure Perception RL-based methods consistently outperform SFT-based methods, with Reason-RFT achieving the best result in 2B model (8% higher than CoT-SFT) while Reason-RFT-Zero performs strongest in 7B model and Reason-RFT remains competitive (16% higher than CoT-SFT), with SFT-based methods showing limited impact especially in the 7B model; (3) Spatial Transformation RL-based methods surpass SFT-based approaches in both 2B and 7B models while significantly outperforming all baseline models, with Reason-RFT (2B) demonstrating remarkable OOD generalization exceeding GPT-4o [26] by 35%. In summary, Reason-RFT surpasses all open-source and proprietary baselines as well as alternative training methods, demonstrating remarkable performance in visual reasoning generalization capabilities.
4.4 Assessment on Training Efficiency
To demonstrate the data efficiency of Reason-RFT during training, we trained all methods on TRANCE and recorded intermediate and validation results, as shown in Fig. 3. Detailed results are provided in the Appendix. Key findings include: (1) Reason-RFT exhibits strong data efficiency. In the ID tasks of the 2B model, Reason-RFT achieves 70% of the performance of Reason-RFT-Zero using only 3% of the training data (1,600 samples), reaching 82.5% with 9% of the data. (2) This efficiency generalizes to OOD tasks. In the 7B model, Reason-RFT achieves over 92% of the performance of Reason-RFT-Zero using just 3% of the training data, demonstrating robust generalization capabilities.
Figure 3: Results of different methods on the Spatial Transformation task across training processes. (a) Evaluation results for 2B model on ID task, (b) Evaluation results for 7B model on ID task, (c) Evaluation results for 2B model on OOD task, (d) Evaluation results for 7B model on OOD task.
4.5 Training Insights
Transient Adaptation Gap This phenomenon refers to the temporary performance drop observed during the initial phase of training process with Reason-RFT-Zero. As the model transitions from directly outputting answers to generating structured reasoning processes, it experiences a brief adaptation difficulty, leading to a sharp performance decline followed by gradual recovery. As shown in Fig. 4 (a), which illustrates the training process of Reason-RFT-Zero on the CLEVR-MATH, both ID and OOD test performances exhibit this sharp drop and recovery within the first 100 steps. We further investigate this phenomenon through a case study, shown in Fig. 4 (b), which demonstrates that forcing the model to output its reasoning process can lead to incorrect answers to originally correct question.
Greedy Reward Stratification This phenomenon describes the model’s tendency during Reason-RFT-Zero training to prioritize easier rewards (e.g., Format Reward) over harder ones (e.g., Accuracy Reward). As shown in Fig. 5, the model’s Reasoning Token Length initially decreases, then gradually increases before stabilizing. This behavior coincides with the Format Reward reaching its initial peak and the Accuracy Reward entering its rapid growth phase. We infer that the model simplifies its outputs early on to quickly adapt to the structured response format, reducing Reasoning Token Length. Once the Format Reward is maximized, the model shifts focus to improving accuracy, increasing Reasoning Token Length.
Figure 4: Illustration of the Transient Adaptation Gap. (a) shows a sharp drop and recovery in both ID and OOD test performances within the early training steps during training on the CLEVR-MATH task. (b) presents a case study of the prediction result on the 10-th step for further investigation.
Figure 5: Analysis of Greedy Reward Stratification. The model’s Reasoning Token Length initially decreases, then gradually increases before stabilizing during Reason-RFT-Zero training. (a) illustrates the curve of Reasoning Token Length, highlighting the Minimum Point. (b) shows the curve of Format Reward, indicating the Maximum Point. (c) displays the curve of Accuracy Reward, marking the Rapid Growth Point. This trend aligns with the Format Reward reaching its initial peak and the Accuracy Reward entering its rapid growth phase.
Figure 6: Analysis of Reasoning Redundancy. (a) illustrates the curves of Reasoning Token Length for Reason-RFT-Zero and Reason-RFT during training on the Structure Perception task, highlighting the significant difference in Reasoning Token Length between these two paradigms in the later steps of training. (b) shows the curves of Accuracy Reward, indicating that the accuracy rates of both paradigms converge and remain nearly identical.
Reasoning Redundancy This phenomenon refers to the significant difference in Reasoning Token Length between models trained under different paradigms. For example, in the Structure Perception task, Reason-RFT and Reason-RFT-Zero achieve comparable final test accuracy, but Reason-RFT exhibits notably longer Reasoning Token Length, as shown in Fig. 6. This occurs because the CoT data used in Reason-RFT’s Reasoning Activation phase, often distilled from stronger models (e.g., GPT-4o), leads the model to learn longer reasoning chains during CoT-SFT. During the Reasoning Reinforcement phase, the model retains these lengthy chains due to the lack of penalties or incentives for response length. In contrast, Reason-RFT-Zero, which lacks CoT data, stabilizes at a shorter Reasoning Token Length through exploration. We hypothesize that Reason-RFT’s longer reasoning chains may be unnecessary for the current task difficulty, introducing redundant computational overhead. Experiments limiting Reasoning Token Length during inference for Reason-RFT-trained models show stable performance until a certain threshold, partially confirming Reasoning Redundancy in specific tasks.
4.6 Exploration on Reward Design
@@
$+$ visual tokens 79.85 58.71 57.98 69.09
@@
Setting $\alpha$ $\beta$ ID OOD AVG
TRANCE TRANCE-L TRANCE-R
@@
Table 5: Results of various mixed CoT activation datasets of Reason-RFT on the Structure Perception task.
Setting CoT Activation Data ID OOD AVG
GeoMath Geometry3k
Baseline GeoMath-only data 59.27 49.25 54.26
(a) Mixed Specific-Domain data 50.61 45.35 48.02
(b) Mixed General-Domain data 42.51 40.25 41.38
Exploration on Format Reward In DeepSeek-R1 [15], the Format Reward mandates the use of <think> tags for structuring reasoning processes in pure-textual tasks. To enhance visual reasoning tasks, we propose integrating visual observations through captioning, extending the format reward with <summary> and <caption> tokens alongside <think>. As shown in Tab. 3, this extension significantly improved the performance of Reason-RFT-Zero, but not Reason-RFT. We attribute this discrepancy to Reason-RFT’s prior exposure to Reasoning Activation, which likely learned the caption-like guidance from a limited CoT reasoning samples, resulting in reduced effectiveness of additional tag guidance. Conversely, Reason-RFT-Zero, being purely RL-based, benefits more from effective tag guidance, highlighting its greater potential for improvement.
@@
Exploration on Accuracy Reward We investigate the Accuracy Reward for the challenging Spatial Transformation task, which requires predicting novel transformation sequences. The reward design in Eq. 7 includes coefficients $\alpha$ and $\beta$ for partial matches, influencing tolerance for incomplete outputs. We test three configurations: (1) $\alpha=0,\beta=0$ (full matches only), (2) $\alpha=0.50,\beta=0.25$ (full and partial matches), and (3) $\alpha=-0.25,\beta=-0.50$ (full matches with partial penalties). Experimental results on 2B and 7B models (Tab. 4) revealed: (1) Rewards for partial matches improved in-domain performance but harmed out-of-domain generalization, suggesting “soft rewards” may limit adaptability; (2) Penalties for partial matches enhanced performance on out-of-domain, indicating “hard rewards” are more effective for serialized tasks.
4.7 Exploration on COT Activation Data
To investigate the impact of differently composed CoT activation data on Reason-RFT, we construct two distinct datasets: a mixed domain-specific dataset, which integrates relevant yet distinct data from in-domain tasks, and a mixed general-domain dataset, encompassing a broader range of visual reasoning tasks (e.g., graph topology, visual puzzles). Using these datasets, we perform Reason-RFT training on Structure Perception task, with the results detailed in Tab. 5. From this, two key points emerge: (1) As the proportion of in-domain training data decreases, the model’s performance on specific tasks declines; (2) Models trained on more diverse visual reasoning domain data may also exhibit a reduction in domain-specific performance.

### v3

@@ 4.1 Experimental Details
Datasets In this paper, we comprehensively evaluate the visual reasoning capabilities of our method by leveraging six existing datasets, enhanced through subtask categorization, error-prone data filtering, and dataset restructuring. Detailed protocols for data filtering and restructuring are provided in Sec. A. Specifically, we define three task categories as follows.
  • Visual Counting is a multimodal reasoning task evaluating the integration of linguistic, visual, and mathematical skills by solving arithmetic problems in 3D block-based scenes. Specifically, we filtered and corrected 35K samples from CLEVR-Math [1] for training and 1K test samples for in-domain (ID) evaluation. To assess generalization under domain-shift (DS), we constructed 1K new samples using 3D assets from Super-CLEVR [51], including two subsets: direct arithmetic (DS-D) and mixed arithmetic (DS-M). Refer to the Appendix Sec. A.1 for details.
  • Structure Perception is a structural reasoning task requiring models to analyze relationships in geometries, imaging structures, chart layouts, and architectural designs. We filtered 4.5K training samples and 820 ID test samples from Geo170K [52] and Math360K [55], along with 800 samples from Geometry3K [85] to evaluate DS adaptability. See the Appendix Sec. A.2.
  • Spatial Transformation is a spatial-visual reasoning task requiring models to infer single- or multi-step transformations by analyzing initial and final states of 3D scenes from multiple perspectives (e.g., center, left, right). We generated 100K samples using TRANCE [56], covering four difficulty levels, and selected 60K for training and 6K for testing through a specific filtering process. For DS evaluation, identical scenes are rendered from left/right viewpoints (DS-L/R) to test perspective-change robustness. Details can be found in the Appendix Sec. A.3.
Evaluation Metrics We use accuracy-rate (Acc) as the primary metric [86]. For numerical answers, correctness is verified by mathematical equivalence to the ground truth. For multiple-choice questions, we perform a string match. For function-type sequences, we use stepwise multi-level evaluation.
Implementation Details We utilize Qwen2-VL-2B and Qwen2-VL-7B [87] as the backbone models for our experiments. Our implementation is built on the open-source frameworks Open-R1 [88] and vLLM [89], ensuring reproducibility and scalability. All experiments were conducted on a cluster of servers, each equipped with $\timesA8$00 GPUs. For further details, see the Appendix Sec. B.
Training Paradigms and Baselines To assess the performance and generalization of different training strategies, we compare: (1) SFT-based methods—ANS-SFT, which fine-tunes on answer generation, and CoT-SFT, which uses supervised learning with CoT reasoning; and (2) RL-based methods—Reason-RFT-Zero, which applies RL without reasoning activation stage, and Reason-RFT, which uses limited CoT data for reasoning activation before RL training. For comprehensive experiments, we use Qwen2-VL-Instruct [87] as the base model (both 2B and 7B variants). In addition, we also select the most advanced open-source models  [90, 91, 92, 93, 94] and the proprietary models  [95, 96] as baselines to evaluate the performance of different paradigms.
4.2 Overall Evaluation of Reason-RFT Framework (RQ1)
To evaluate Reason-RFT, we evaluate Reason-RFT using 2B- and 7B-parameter models on three visual reasoning tasks. The results are summarized as follows.
Strong reasoning performance across ID tasks. As shown in Tab. 1, Reason-RFT achieves performance comparable to or better than both SFT- and RL-based methods across all tasks. In visual counting, Reason-RFT-Zero achieves the best performance among all models in the 7B setting. In structure perception, Reason-RFT outperforms most open-source and proprietary baselines in the 7B setting and remains competitive with top models such as InternVL-2.5-8B [92]. In spatial transformation, Reason-RFT matches or exceeds SFT-based methods while consistently outperforming all baselines. These results demonstrate that Reason-RFT effectively integrates the strengths of both SFT and RL in structured reasoning tasks.
Superior generalization under DS. Under DS settings, Reason-RFT shows substantial gains over both traditional baselines and alternative training paradigms. In visual counting, it outperforms ANS-SFT by 10.95% (2B) and 13.93% (7B). In structure perception, Reason-RFT achieves the highest performance in the 2B model, with an 6.93% gain over CoT-SFT, and remains highly competitive in the 7B model. Most notably, in the spatial transformation task, the 2B Reason-RFT model surpasses GPT-4o [95] by 31.7%, showcasing remarkable generalization under DS.
Figure 3: Results of different methods and model sizes on spatial transformation task across training.
High training efficiency. To evaluate data efficiency during training, we train all methods on the spatial transformation task and monitor intermediate ID and DS performance (Fig. 3). Additional results for visual counting and structure perception are provided in Appendix Sec. C. In the 2B model, Reason-RFT achieves 70% of the final performance of Reason-RFT-Zero using only 3% of the training data (1,600 samples), and reaches 82.5% with 9% of the data. In the 7B model, it achieves over 92% of Reason-RFT-Zero’s performance using just 3% of the data. These results confirm that Reason-RFT exhibits strong training efficiency in both ID and DS scenarios. The data-efficient nature of Reason-RFT renders it particularly effective for few-shot learning, offering significant potential for applications where labeled data is limited.
@@ 4.2 Overall Evaluation of Reason-RFT Framework (RQ1)
Table 1: Results on three visual reasoning tasks. The best results among different training paradigms are highlighted in bold, while the second-best results are underlined. “ID” denotes in-domain test data, and “DS” denotes domain-shift test data.
@@
ID DS-D DS-M AVG ID DS AVG ID DS-L DS-R AVG
@@
Qwen2.5-VL-3B-Inst. [90] 75.90 50.93 4.40 43.74 36.75 37.44 37.09 8.57 8.26 8.31 8.42
Phi-3.5-Vision-4B-Inst. [91] 21.40 18.27 6.00 15.22 36.83 50.25 43.54 7.42 2.45 4.02 5.33
Qwen2.5-VL-7B-Inst. [90] 74.60 46.00 2.80 41.13 44.00 45.61 44.80 19.63 13.12 13.42 16.45
InternVL-2.5-8B [92] 93.50 46.13 2.80 47.48 63.00 47.32 51.60 7.19 6.62 6.63 6.91
Llama-3.2-11B-Vision [93] 10.30 9.60 9.20 9.70 13.75 20.85 17.30 8.22 8.40 9.03 8.47
Pixtral-12B [94] 42.60 25.33 15.60 27.84 30.38 36.09 33.23 7.35 5.03 5.22 6.42
@@
Zero-Shot 82.40 42.67 0.00 41.69 25.86 20.63 23.25 3.78 4.60 4.67 4.35
++ ANS-SFT 96.20 51.07 5.20 50.82 51.34 22.50 36.92 77.39 49.24 50.33 58.99
++ CoT-SFT 85.50 49.73 36.80 57.34 43.05 25.25 34.15 64.37 43.19 42.86 50.14
++ Reason-RFT-Zero 98.40 58.00 5.20 53.87 47.68 32.50 40.09 42.13 34.07 33.41 33.74
++ Reason-RFT 96.80 60.00 28.40 61.77 49.03 33.13 41.08 74.61 64.05 64.08 67.58
@@
Zero-Shot 98.60 54.53 4.80 52.64 43.30 43.88 43.59 13.53 12.72 12.78 13.01
++ ANS-SFT 95.00 42.53 8.00 48.51 51.34 25.38 38.36 82.19 54.29 54.83 63.77
++ CoT-SFT 87.30 45.33 33.60 55.41 50.49 33.00 41.75 81.31 47.90 47.80 59.00
++ Reason-RFT-Zero 99.40 63.60 21.20 61.40 55.00 54.75 54.88 67.67 57.20 56.15 60.34
++ Reason-RFT 95.60 56.13 35.60 62.44 59.27 49.25 54.26 79.97 59.36 58.61 65.98
4.3 Effect of STAGE 1 on Initialization (RQ2)
To investigate the role of CoT-SFT in initialization, we compare four baselines across three tasks. The results in Tab. 1 reveal the following: (1) Consistent performance gains from CoT-SFT. Across all three tasks and both 2B and 7B model scales, Reason-RFT consistently outperforms Reason-RFT-Zero following stage 1 reasoning activation. This improvement is particularly notable when the model is small and the task involves complex output structures. For example, in the spatial transformation task—which requires function-like serialized outputs—the 2B Reason-RFT model surpasses Reason-RFT-Zero by 33.84%. (2) Smaller models benefit more from CoT-SFT priors. In the visual counting task under the DS-M setting, the 2B model with CoT-SFT outperforms Reason-RFT-Zero by 31.6%. Although the gap narrows in the 7B model, CoT-SFT still yields substantial gains. This indicates that pure RL-based methods struggle to adapt from direct arithmetic to mixed arithmetic reasoning under DS, whereas CoT-SFT provides effective inductive priors for such adaptation. Moreover, under the same amount of CoT-SFT data, the 2B model still underperforms its 7B counterpart, highlighting the increased reliance of smaller models on CoT-SFT for acquiring reasoning capabilities.
Figure 4: Results of DS v.s. ID on spatial transformation task.
4.4 Effect of STAGE 2 on Generalization (RQ3)
To evaluate the impact of RL in stage 2, we compare the generalization performance of CoT-SFT and Reason-RFT across three visual reasoning tasks under DS. As shown in Tab. 1 and Fig. 4, the results reveal the following: Reinforcement fine-tuning improves generalization beyond CoT-SFT. Across all domain-shift settings, Reason-RFT consistently outperforms CoT-SFT, demonstrating that reinforcement learning significantly enhances model robustness. For instance, in the visual counting task, the 7B Reason-RFT achieves a combined DS-D and DS-M score 12.8% higher than CoT-SFT. The improvement is even more pronounced in structure-sensitive tasks such as spatial transformation, where the 2B Reason-RFT exceeds CoT-SFT by 21.04% on average across DS-L and DS-R. These results indicate that CoT-SFT alone yields limited generalization, while reinforcement fine-tuning enables better adaptation to compositional and layout-dependent variations.
4.5 Exploration on Reward Design (RQ4)
@@
++ visual tokens 79.85 58.71 57.98 69.09
Format Reward. In DeepSeek-R1 [39], the format reward enforces the use of <think> and <answer> tokens to structure reasoning in textual tasks. To better support visual reasoning, we extend this with <summary> and <caption> tokens to incorporate visual observations via caption-style prompts. As shown in Tab. 2, this improves Reason-RFT-Zero but has limited effect on Reason-RFT. We attribute this to Reason-RFT’s prior CoT supervision, which likely helps it internalize caption-like structures in stage 1, reducing the benefit of explicit tags. In contrast, Reason-RFT-Zero benefits more from such structural cues, indicating greater sensitivity to format-level guidance.
@@
Setting $\alpha$ $\beta$ ID DS-L DS-R AVG
@@
Accuracy Reward. We explore accuracy reward design in the spatial transformation task, which requires predicting transformation sequences in a structured format. The formulation in Eq. 7 introduces coefficients $\alpha$ and $\beta$ to control tolerance for partial matches. We test three settings: (1) $\alpha=0,\beta=0$ (exact match only), (2) $\alpha=0.50,\beta=0.25$ (partial credit), and (3) $\alpha=-0.25,\beta=-0.50$ (penalized partial matches). Results on 2B and 7B models (Tab. 3) show that: (1) partial credit improves ID performance but harms generalization, suggesting “soft rewards” reduce robustness; (2) penalizing partial matches improves generalization under domain shift, indicating “hard rewards” better support serialized reasoning.
Figure 5: Analysis of Greedy Reward Stratification. The model’s reasoning token length first decreases, then gradually rises and stabilizes during Reason-RFT-Zero training. The peak of the format reward coincides with the accuracy reward’s rapid growth phase.
Figure 6: Illustration of the Transient Adaptation Gap. (a) shows a sharp drop and recovery in both ID and DS test performances within the early training steps. (b) shows a case study of the prediction result on early step.
Figure 7: Analysis of Reasoning Redundancy. (a) shows the reasoning token length curves for Reason-RFT-Zero and Reason-RFT during training. (b) displays their accuracy reward, with both paradigms converging to similar accuracy rate.
4.6 Training Dynamics and Reasoning Behavior Analysis (RQ5)
Greedy Reward Stratification. This phenomenon captures the model’s tendency, particularly in Reason-RFT-Zero, to prioritize easier-to-optimize rewards (e.g., format reward) before addressing more challenging objectives (e.g., accuracy reward). As shown in Fig. 5, the reasoning token length initially drops, then gradually increases and stabilizes. This dynamic correlates with the format reward quickly reaching a plateau, followed by a sharp rise in the accuracy reward. We infer that the model initially simplifies its outputs to adapt rapidly to structured format expectations, and only later allocates learning capacity to improve reasoning correctness.
Transient Adaptation Gap. This refers to a temporary performance degradation observed in the early training phase of Reason-RFT-Zero. When the model shifts from directly predicting answers to producing structured reasoning traces, it experiences a brief adaptation bottleneck—marked by a sharp decline and gradual recovery in accuracy. Fig. 7 (a) illustrates this drop within the first 100 steps on the visual counting task under both ID and DS settings. A case analysis in Fig. 7 (b) further reveals that forcing structured reasoning prematurely may lead to incorrect outputs, highlighting the sensitivity of early-stage training to reasoning format constraints.
Reasoning Redundancy. This phenomenon concerns the discrepancy in reasoning token length between models trained with and without CoT activation. In the structure perception task, Reason-RFT and Reason-RFT-Zero attain similar accuracy, yet the former generates significantly longer reasoning traces (Fig. 7). This likely stems from Reason-RFT’s use of CoT data distilled from advanced models (e.g., GPT-4o), which encourages verbose reasoning during stage 1. In the absence of penalties or length control in reinforcement fine-tuning, such verbosity persists. By contrast, Reason-RFT-Zero converges to more concise reasoning through reward-driven exploration. We hypothesize that these longer chains in Reason-RFT may introduce unnecessary computational overhead or reflect overthinking relative to task complexity.